### v2

@@ Abstract
In an era of increasing computational capabilities and growing environmental consciousness, organizations face a critical challenge in balancing the accuracy of forecasting models with computational efficiency and sustainability. Global forecasting models, lowering the computational time, have gained significant attention over the years. However, the common practice of retraining these models with new observations raises important questions about the costs of forecasting. Using ten different machine learning and deep learning models, we analyzed various retraining scenarios, ranging from continuous updates to no retraining at all, across two large retail datasets. We showed that less frequent retraining strategies maintain the forecast accuracy while reducing the computational costs, providing a more sustainable approach to large-scale forecasting. We also found that machine learning models are a marginally better choice to reduce the costs of forecasting when coupled with less frequent model retraining strategies as the frequency of the data increases. Our findings challenge the conventional belief that frequent retraining is essential for maintaining forecasting accuracy. Instead, periodic retraining offers a good balance between predictive performance and efficiency, both in the case of point and probabilistic forecasting. These insights provide actionable guidelines for organizations seeking to optimize forecasting pipelines while reducing costs and energy consumption.
@@
Time series, Demand forecasting, Forecasting competitions, Cross-learning, Global models, Machine learning, Deep learning, Green AI, Conformal predictions
@@ 1 Introduction
Forecasting plays a critical role in decision-making processes across industries, from supply chain management to energy demand planning. Traditionally, time series forecasting has relied on a Local Modeling approach (LM), meaning that separate models 111 We use the words model and method interchangeably, even if the two have distinct statistical meanings (Svetunkov & Boylan, 2023). are trained for each series in isolation. However, recent advances have demonstrated the potential of Global forecasting Models (GM), a paradigm also known as cross-learning, which in turn consists of fitting a single forecasting model on the whole data set (Januschowski et al., 2020). This novel approach allows to leverage information across multiple time series to improve accuracy and generalization by employing machine learning or deep learning architectures 222 We distinguish between traditional machine learning models, like linear regression or tree-based methods, and deep learning models, those that are based on some sort of neural network architecture. . Global modeling has affirmed itself as one of the most relevant innovations in time series forecasting of recent years (Semenoglou et al., 2021), and its success is not only due to the demonstrated forecast accuracy in most time series competitions, but also to the particular computational benefit it provides. Indeed, while the local approach to time series forecasting is potentially beneficial for forecast accuracy, it comes with substantial computational overhead. Moreover, since the forecasts are usually produced through some cloud computing service that relies on pay-as-you-go pricing (Fotios Petropoulos & Spiliotis, 2024), higher computing time and resources directly translate into higher costs of forecasting for organizations.
@@
The retail industry particularly exemplifies the complexity of modern forecasting challenges. Retailers must predict demand for thousands or even millions of products across hundreds of stores while accounting for seasonality, promotions, pricing strategies, and external factors such as weather and local events. Accurate demand forecasts directly impact inventory decisions: overestimation leads to excess inventory and waste, particularly critical for perishable goods, while underestimation results in stockouts and direct money loss for companies (Fildes et al., 2022). Traditional approaches involving separate models for each product-store combination have proven increasingly inadequate, especially in contexts where the number of series (SKUs) to forecast is high, leading to the adoption of global forecasting models that can learn patterns across the entire product hierarchy and can drastically reduce the costs of producing forecasts 333 Most of the cost of producing forecasts is due to the algorithm’s training time. The global modeling approach cuts it by training the model only once on the whole data set. .
@@
Nevertheless, despite the growing adoption of global models, it is still common practice to update the forecasting models when new observations are available, often motivated by the assumption that continuous updates lead to better adaptability to changing patterns and more accurate predictions. Unlike local models, global models benefit from learning shared dynamics across multiple time series, potentially leading to more robust representations that are less sensitive to frequent retraining. In the context of local models, the seminal work of Spiliotis & Petropoulos (2024) demonstrated that less frequent model updates do not harm forecasting accuracy. However, in the context of global models, the effects of retraining are not well understood. Is continuous retraining necessary to maintain forecast performance, or can global models remain effective without frequent updates? This aspect remains largely unexplored in the forecasting literature, and investigating the impact of retraining is crucial for both methodological and practical reasons. From a methodological perspective, understanding whether global models degrade in performance without frequent retraining provides insights into their stability and adaptability. If global models can maintain strong predictive accuracy with less frequent updates, it would challenge conventional wisdom on the necessity of continuous retraining in forecasting. Moreover, a stable forecasting model provides great benefit to the whole decision process the forecasts are used for. From a practical perspective, frequent retraining has significant computational and environmental costs. Training large-scale forecasting models requires substantial computational power, contributing to energy consumption and carbon emissions (Schwartz et al., 2020). Indeed, the energy consumption of model training extends beyond direct computational costs. In recent years, the environmental impact of machine learning models has become a growing concern. Data centers running these models contribute significantly to global carbon emissions, with estimates suggesting that training a single large deep learning model can emit as much carbon as five cars over their lifetimes (Strubell et al., 2019). The frequency of model retraining multiplies this impact, particularly in large organizations handling millions of series, emphasizing the need for more computationally efficient training practices. The implication is straightforward: reducing the retraining frequency of forecasting models could contribute substantially to sustainability by lowering energy consumption, possibly without harming accuracy performance.
@@ 1.1 Research Question
We aim to extend the results presented in the influential work of Spiliotis & Petropoulos (2024), addressing the question ”Is frequent retraining necessary in the context of global forecasting models?”. Specifically, we study whether skipping retraining when new observations are available harms the forecasting performance of global models. To answer this question, we use the two most recent and comprehensive retail forecasting datasets, namely the M5 and the VN1 competitions’ data.
@@ 2 Related works
The literature on cross-learning approach has evolved significantly in recent years. Nowadays, most of the works related to time series forecasting include at least some benchmark comparison with global models, demonstrating their relevance in the field. Semenoglou et al. (2021) extensively showed their accuracy on the M4 competition dataset, Hewamalage et al. (2022) evaluated the conditions when global forecasting models are competitive, and Montero-Manso & Hyndman (2021) and Montero-Manso (2023) theoretically demonstrated that GM are at least as accurate as local models with less complexity and without any assumption on the similarity of the data. GM emerged as the most accurate approach in many forecasting areas, such as gas consumption (Gaweł & Paliński, 2024), electricity demand (Buonanno et al., 2022), water demand (Groß & Hans, 2024), crop production (Ibañez & Monterola, 2023), and retail demand (Spiliotis et al. (2022), Bandara et al. (2019), Juan R Trapero & Fildes (2015)). Moreover, GM effectiveness shined during the M5 competition (Makridakis et al., 2022a), where tree-based models leveraging cross-learning were among almost all the top forecasting solutions (Januschowski et al., 2022). Several techniques like clustering (Godahewa et al. (2021a), Bandara et al. (2020)), and data augmentation (Bandara et al., 2021) have been tested to further increase the performance of GM. Furthermore, new machine learning (Godahewa et al., 2023) and deep learning (Oreshkin et al., 2020) architectures specifically designed for cross-learning have been developed. Finally, a novel area of research is emerging that tries to improve the ability of GM to capture local patterns (Sen et al., 2019), and their explainability (Rajapaksha et al., 2023).
@@
From a forecasting evaluation perspective, most of the literature on GM is focused on point prediction accuracy, possibly because most machine learning and deep learning methods do not directly output probabilistic forecasts (Makridakis et al., 2022c). Nevertheless, in many forecasting contexts (like supply chain) it is very important to be able to produce and evaluate predictions in a probabilistic way (being intervals, quantiles, or density based) (Fildes et al., 2022). Among others, Vovk et al. (2005) introduced a novel tool for uncertainty quantification under any machine learning model, namely Conformal Inference, that can also be applied in time series forecasting experiments (Stankeviciute et al., 2021).
@@
In the context of model retraining and updating strategies, instead, Spiliotis & Petropoulos (2024) is the main work related to time series forecasting. The authors explored extensively the effects of different retraining scenarios and different forms of model parameter updates on the model performance, although they focused on the exponential smoothing family of models following the traditional local approach. Huber & Stuckenschmidt (2020) briefly discussed retraining in the context of retail demand, but with few models and retraining scenarios, and on a proprietary daily dataset only. Despite the findings being promising, there has been little direct exploration of whether global models specifically require frequent retraining or if they can retain competitive accuracy with less frequent updates. However, this topic is under consideration of the broader machine learning community (Getzner et al., 2023), advocating for Green AI (Schwartz et al., 2020).
@@ 3.1 Datasets
For the experiment, we used two retail forecasting datasets: the M5 and the VN1 competition datasets. The M5 competition was organized by Spyros Makridakis and his colleagues as part of the M-competitions series, which aimed to compare different forecasting methods in the context of retail demand (Makridakis et al., 2022b). The M5 (Howard & Makridakis, 2020) is a well-known and well-studied dataset containing 30,490 SKU $\times$ Store daily time series of Walmart’s unit sales of products. It covers the sales of three categories of products (Food, Hobbies, and Household) sold into ten different stores located in three US states (California, Texas, and Wisconsin). The time period spans from 2011 to 2016. The time series are highly intermittent and are hierarchically organized, allowing forecasting at multiple levels, such as individual products, product categories, stores, and States. Exogenous information that can influence sales, such as product prices, promotions, and special events (e.g., holidays) are also available. The VN1 Forecasting - Accuracy Challenge competition was jointly organized by Flieber, Syrup Tech, and SupChains in October 2024, and it is the first of its series (Vandeput, 2024). The dataset contains weekly sales of 15,053 products sold from 2020 until 2024 from e-vendors. In particular, they were mostly online from the US, and the products were directly sold to the final consumers. Contrary to the M5 dataset, where all the products were sold by a single retailer (Walmart) and from just a few stores, the VN1 dataset includes product sales of 328 warehouses from 46 different retailers. As far as we know, we are among the first to test forecasting models on this data. These sets of data are the most recent and comprehensive time series datasets related to retail demand, allowing for a good generalization of the results, in particular in the context of demand forecasting.
Table 1: Characteristics of the different datasets used in the experiments. The number of series is based on the minimum number of observations per series.
Dataset Frequency N. Series Min Obs per Series
M5 Daily (7) 28.298 730
VN1 Weekly (52) 15.053 157
@@ 3.2 Forecasting models
Let us define $\mathcal{Y}$ as the set of all available time series in a dataset, such that $Y_{i}$ represents a single component, and let $\mathcal{F}$ be the set of possible predictive functions, so that $F$ corresponds to a single model 444 In our setting, $F$ can be any predictive model in machine learning or deep learning framework. We do not consider classical statistical forecasting methods, like ARIMA or Exponential Smoothing, since in their common formulations are purely local models. . Without loss of generality, we can assume that all the necessary information for prediction are contained in $\mathcal{Y}$. Under the local approach, predictions for the forecast horizon $h$ are obtained training a model for each time series in the data set, implying that each time series has its own model, defined by its own parameters values.
@@ 3.2.1 Machine learning models
Linear Regression (LR) is a classical statistical model that assumes a linear relationship between input features and the target variable. Despite its simplicity, LR can be effective for time series forecasting when combined with appropriate feature engineering. The Pooled Regression is considered a solid benchmark for global model performance evaluation (Montero-Manso & Hyndman, 2021; Godahewa et al., 2021b), and it has also proven to be quite effective (Bandara et al., 2022).
Random Forest (RF) is an ensemble learning method based on regression trees. It constructs multiple trees during training and aggregates their predictions through averaging. RF excels in capturing non-linear patterns and interactions between variables, and it is robust to overfitting, making it a strong tree-based model for demand forecasting (Breiman, 2001). It was the method used by Amazon until 2015 to forecast e-commerce products demand (Januschowski et al., 2022) 555 For the cost of computation, the Random Forest model is tested only on the VN1 weekly dataset. .
Extreme Gradient Boosting (XGBoost) is a gradient-boosted decision tree algorithm designed for speed and performance. Iteratively optimizing an objective function combines weak learners’ predictions to form a solid predictive model (Chen & Guestrin, 2016). XGBoost incorporates regularization techniques to prevent overfitting, making it particularly effective for datasets with complex relationships (Shwartz-Ziv & Armon, 2022).
Light Gradient Boosting Machine (LGBM) is another gradient-boosted tree algorithm that emphasizes efficiency and scalability. It uses a histogram-based approach to split features and a leaf-wise growth strategy to reduce computation time. LGBM is particularly well-suited for large datasets and high-dimensional data, often outperforming other boosting algorithms in terms of speed and accuracy (Ke et al., 2017). In demand forecasting, it is among the top solutions of all the major competitions (Favorita, Rossmann, M5, and VN1) (Makridakis et al., 2022a, c; In & Jung, 2022; Lainder & Wolfinger, 2022).
@@ 3.2.2 Deep learning models
Recurrent Neural Networks (RNN) are designed to model sequential data by maintaining a hidden state that captures temporal dependencies. Standard RNNs, however, suffer from vanishing gradient problems, limiting their ability to learn long-term dependencies (Cho et al., 2014). Variants such as Long Short-Term Memory (LSTM) networks and Gated Recurrent Units (GRUs) address this limitation and are widely used for time series forecasting (Hochreiter & Schmidhuber, 1997) 666 In our work, we used only the LSTM version, but, for the cost of computation, it was trained only on the VN1 weekly dataset. .
@@
Neural Basis Expansion Analysis for Time Series (NBEATS) is a deep learning model specifically designed for time series forecasting. It employs a sequence of fully connected layers organized into blocks. Each block outputs both trend and seasonal components, enabling the model to decompose and predict time series effectively. The network has an interpretable configuration that sequentially projects the signal into polynomials and harmonic basis to learn trend and seasonality components (Oreshkin et al., 2020). The Neural Basis Expansion Analysis with Exogenous (NBEATSx), allows for the inclusion of exogenous temporal variables available at the time of the prediction (Olivares et al., 2023). This method has shown state-of-the-art performance on several benchmark datasets and competitions (Anderer & Li, 2022).
Neural Hierarchical Interpolation for Time Series (NHITS) builds upon the success of NBEATS by incorporating hierarchical interpolation mechanisms to better capture time hierarchies in time series data. Multi-rate input pooling, hierarchical interpolation, and backcast residual connections together induce the specialization of the additive predictions in different signal bands, reducing memory and computational time, thus improving the architecture’s parsimony and accuracy (Challu et al., 2022).

### v3

@@ Abstract
In an era of increasing computational capabilities and growing environmental consciousness, organizations face a critical challenge in balancing the accuracy of forecasting models with computational efficiency and sustainability. Global forecasting models, lowering the computational time, have gained significant attention over the years. However, the common practice of retraining these models with new observations raises important questions about the costs of forecasting. Using ten different machine learning and deep learning models, we analyzed various retraining scenarios, ranging from continuous updates to no retraining at all, across two large retail demand datasets. We showed that less frequent retraining strategies maintain the forecast accuracy while reducing the computational costs, providing a more sustainable approach to large-scale forecasting. We also found that machine learning models are a marginally better choice to reduce the costs of forecasting when coupled with less frequent model retraining strategies as the frequency of the data increases. Our findings challenge the conventional belief that frequent retraining is essential for maintaining forecasting accuracy. Instead, periodic retraining offers a good balance between predictive performance and efficiency, both in the case of point and probabilistic forecasting. These insights provide actionable guidelines for organizations seeking to optimize forecasting pipelines while reducing costs and energy consumption.
@@
Forecasting , Time series , Machine learning , Deep learning , Green AI
journal: Machine Learning with Applications
@@ 1 Introduction
Forecasting plays a critical role in decision-making processes across industries, from supply chain management to energy demand planning. Traditionally, time series forecasting has relied on a Local Modeling approach (LM), meaning that separate models are trained for each series in isolation. However, recent advances have demonstrated the potential of Global forecasting Models (GM), a paradigm also known as cross-learning, which in turn consists of fitting a single forecasting model on the whole data set (Januschowski et al., 2020). This novel approach allows to leverage information across multiple time series to improve accuracy and generalization by employing machine learning or deep learning architectures. Global modeling has affirmed itself as one of the most relevant innovations in time series forecasting of recent years (Semenoglou et al., 2021), and its success is not only due to the demonstrated forecast accuracy in most time series competitions, but also to the particular computational benefit it provides. Indeed, while the local approach to time series forecasting is potentially beneficial for forecast accuracy, it comes with substantial computational overhead. Moreover, since the forecasts are usually produced through some cloud computing service that relies on pay-as-you-go pricing (Petropoulos and Spiliotis, 2024), higher computing time and resources directly translate into higher costs of forecasting for organizations.
@@
The retail industry particularly exemplifies the complexity of modern forecasting challenges. Retailers must predict demand for thousands or even millions of products across hundreds of stores while accounting for seasonality, promotions, pricing strategies, and external factors such as weather and local events. Accurate demand forecasts directly impact inventory decisions: overestimation leads to excess inventory and waste, particularly critical for perishable goods, while underestimation results in stockouts and direct money loss for companies (Fildes et al., 2022). Traditional approaches involving separate models for each product-store combination have proven increasingly inadequate, especially in contexts where the number of series (SKUs) to forecast is high, leading to the adoption of global forecasting models that can learn patterns across the entire product hierarchy and can drastically reduce the costs of producing forecasts.
@@
Nevertheless, despite the growing adoption of global models, it is still common practice to update the forecasting models when new observations are available, often motivated by the assumption that continuous updates lead to better adaptability to changing patterns and more accurate predictions. Unlike local models, global models benefit from learning shared dynamics across multiple time series, potentially leading to more robust representations that are less sensitive to frequent retraining. In the context of local models, the seminal work of Spiliotis and Petropoulos (2024) demonstrated that less frequent model updates do not harm forecasting accuracy. However, in the context of global models, the effects of retraining are not well understood. Is continuous retraining necessary to maintain forecast performance, or can global models remain effective without frequent updates? This aspect remains largely unexplored in the forecasting literature, and investigating the impact of retraining is crucial for both methodological and practical reasons. From a methodological perspective, understanding whether global models degrade in performance without frequent retraining provides insights into their stability and adaptability. If global models can maintain strong predictive accuracy with less frequent updates, it would challenge conventional wisdom on the necessity of continuous retraining in forecasting. Moreover, a stable forecasting model provides great benefit to the whole decision process the forecasts are used for. From a practical perspective, frequent retraining has significant computational and environmental costs. Training large-scale forecasting models requires substantial computational power, contributing to energy consumption and carbon emissions (Schwartz et al., 2020). Indeed, the energy consumption of model training extends beyond direct computational costs. In recent years, the environmental impact of machine learning models has become a growing concern. Data centers running these models contribute significantly to global carbon emissions, with estimates suggesting that training a single large deep learning model can emit as much carbon as five cars over their lifetimes (Strubell et al., 2019). The frequency of model retraining multiplies this impact, particularly in large organizations handling millions of series, emphasizing the need for more computationally efficient training practices. The implication is straightforward: reducing the retraining frequency of forecasting models could contribute substantially to sustainability by lowering energy consumption, possibly without harming accuracy performance.
@@ 1.1 Research Question
We aim to extend the results presented in the influential work of Spiliotis and Petropoulos (2024), addressing the question ”Is frequent retraining necessary in the context of retail demand global forecasting models?”. Specifically, we study whether skipping retraining when new observations are available harms the forecasting performance of global models. To answer this question, we use the two most recent and comprehensive retail forecasting datasets, namely the M5 and the VN1 competitions’ data.
@@ 2 Related works
The literature on cross-learning approach has evolved significantly in recent years. Nowadays, most of the works related to time series forecasting include at least some benchmark comparison with global models, demonstrating their relevance in the field. Semenoglou et al. (2021) extensively showed their accuracy on the M4 competition dataset, Hewamalage et al. (2022) evaluated the conditions when global forecasting models are competitive, and Montero-Manso and Hyndman (2021) and Montero-Manso (2023) theoretically demonstrated that GM are at least as accurate as local models with less complexity and without any assumption on the similarity of the data. GM emerged as the most accurate approach in many forecasting areas, such as gas consumption (Gaweł and Paliński, 2024), electricity demand (Buonanno et al., 2022), water demand (Groß and Hans, 2024), crop production (Ibañez and Monterola, 2023), and retail demand (Spiliotis et al. (2022), Bandara et al. (2019), Juan R Trapero and Fildes (2015)). Moreover, GM effectiveness shined during the M5 competition (Makridakis et al., 2022b), where tree-based models leveraging cross-learning were among almost all the top forecasting solutions (Januschowski et al., 2022). Several techniques like clustering (Godahewa et al. (2021a), Bandara et al. (2020)), and data augmentation (Bandara et al., 2021) have been tested to further increase the performance of GM. Furthermore, new machine learning (Godahewa et al., 2023) and deep learning (Oreshkin et al., 2020) architectures specifically designed for cross-learning have been developed. Finally, a novel area of research is emerging that tries to improve the ability of GM to capture local patterns (Sen et al., 2019), and their explainability (Rajapaksha et al., 2023).
@@
From a forecasting evaluation perspective, most of the literature on GM is focused on point prediction accuracy, possibly because most machine learning and deep learning methods do not directly output probabilistic forecasts (Makridakis et al., 2022a). Nevertheless, in many forecasting contexts (like supply chain) it is very important to be able to produce and evaluate predictions in a probabilistic way, being intervals, quantiles, or density based (Fildes et al., 2022).
@@
In the context of model retraining and updating strategies, instead, Spiliotis and Petropoulos (2024) is the main work related to time series forecasting. The authors explored extensively the effects of different retraining scenarios and different forms of model parameter updates on the model performance, although they focused on the exponential smoothing family of models following the traditional local approach. Huber and Stuckenschmidt (2020) briefly discussed retraining in the context of retail demand, but with few models and retraining scenarios, and on a proprietary daily dataset only. Despite the findings being promising, there has been little direct exploration of whether global models specifically require frequent retraining or if they can retain competitive accuracy with less frequent updates. However, this topic is under consideration of the broader machine learning community (Getzner et al., 2023), advocating for Green AI (Schwartz et al., 2020).
@@ 3.1 Datasets
For the experiment, we used two retail forecasting datasets: the M5 and the VN1 competition datasets. The M5 competition was organized by Spyros Makridakis and his colleagues as part of the M-competitions series, which aimed to compare different forecasting methods in the context of retail demand (Makridakis et al., 2022c). The M5 (Howard and Makridakis, 2020) is a well-known and well-studied dataset containing 30,490 SKU $\times$ Store daily time series of Walmart’s unit sales of products. It covers the sales of three categories of products (Food, Hobbies, and Household) sold into ten different stores located in three US states (California, Texas, and Wisconsin). The time period spans from 2011 to 2016. The time series are highly intermittent and are hierarchically organized, allowing forecasting at multiple levels, such as individual products, product categories, stores, and States. Exogenous information that can influence sales, such as product prices, promotions, and special events (e.g., holidays) are also available. The VN1 Forecasting - Accuracy Challenge competition was jointly organized by Flieber, Syrup Tech, and SupChains in October 2024, and it is the first of its series (Vandeput, 2024). The dataset contains weekly sales of 15,053 products sold from 2020 until 2024 from e-vendors. In particular, they were mostly online from the US, and the products were directly sold to the final consumers. Contrary to the M5 dataset, where all the products were sold by a single retailer (Walmart) and from just a few stores, the VN1 dataset includes product sales of 328 warehouses from 46 different retailers. As far as we know, we are among the first to test forecasting models on this data. These sets of data are the most recent and comprehensive time series datasets related to retail demand, allowing for a good generalization of the results, in particular in the context of demand forecasting.
@@ 3.2 Forecasting models
Let us define $\mathcal{Y}$ as the set of all available time series in a dataset, such that $Y_{i}$ represents a single component, and let $\mathcal{F}$ be the set of possible predictive functions, so that $F$ corresponds to a single model. Without loss of generality, we can assume that all the necessary information for prediction are contained in $\mathcal{Y}$. Under the local approach, predictions for the forecast horizon $h$ are obtained training a model for each time series in the data set, implying that each time series has its own model, defined by its own parameters values.
@@ 3.2.1 Machine learning models
Linear Regression (LR) is a classical statistical model that assumes a linear relationship between input features and the target variable. Despite its simplicity, LR can be effective for time series forecasting when combined with appropriate feature engineering. The Pooled Regression is considered a solid benchmark for global model performance evaluation (Montero-Manso and Hyndman, 2021; Godahewa et al., 2021b), and it has also proven to be quite effective (Bandara et al., 2022).
Random Forest (RF) is an ensemble learning method based on regression trees. It constructs multiple trees during training and aggregates their predictions through averaging. RF excels in capturing non-linear patterns and interactions between variables, and it is robust to overfitting, making it a strong tree-based model for demand forecasting (Breiman, 2001). It was the method used by Amazon until 2015 to forecast e-commerce products demand (Januschowski et al., 2022)  [^1] For the cost of computation, the Random Forest model is tested only on the VN1 weekly dataset. .
Extreme Gradient Boosting (XGBoost) is a gradient-boosted decision tree algorithm designed for speed and performance. Iteratively optimizing an objective function combines weak learners’ predictions to form a solid predictive model (Chen and Guestrin, 2016). XGBoost incorporates regularization techniques to prevent overfitting, making it particularly effective for datasets with complex relationships (Shwartz-Ziv and Armon, 2022).
Light Gradient Boosting Machine (LGBM) is another gradient-boosted tree algorithm that emphasizes efficiency and scalability. It uses a histogram-based approach to split features and a leaf-wise growth strategy to reduce computation time. LGBM is particularly well-suited for large datasets and high-dimensional data, often outperforming other boosting algorithms in terms of speed and accuracy (Ke et al., 2017). In demand forecasting, it is among the top solutions of all the major competitions (Favorita, Rossmann, M5, and VN1) (Makridakis et al., 2022b, a; In and Jung, 2022; Lainder and Wolfinger, 2022).
@@ 3.2.2 Deep learning models
Recurrent Neural Networks (RNN) are designed to model sequential data by maintaining a hidden state that captures temporal dependencies. Standard RNNs, however, suffer from vanishing gradient problems, limiting their ability to learn long-term dependencies (Cho et al., 2014). Variants such as Long Short-Term Memory (LSTM) networks and Gated Recurrent Units (GRUs) address this limitation and are widely used for time series forecasting (Hochreiter and Schmidhuber, 1997)  [^2] We used only the LSTM version, but, for the cost of computation, it was trained only on the VN1 dataset. .
@@
Neural Basis Expansion Analysis for Time Series (NBEATS) is a deep learning model specifically designed for time series forecasting. It employs a sequence of fully connected layers organized into blocks. Each block outputs both trend and seasonal components, enabling the model to decompose and predict time series effectively. The network has an interpretable configuration that sequentially projects the signal into polynomials and harmonic basis to learn trend and seasonality components (Oreshkin et al., 2020). The Neural Basis Expansion Analysis with Exogenous (NBEATSx), allows for the inclusion of exogenous temporal variables available at the time of the prediction (Olivares et al., 2023). This method has shown state-of-the-art performance on several benchmark datasets and competitions (Anderer and Li, 2022).
Neural Hierarchical Interpolation for Time Series (NHITS) builds upon the success of NBEATS by incorporating hierarchical interpolation mechanisms to better capture time hierarchies in time series data. Multi-rate input pooling, hierarchical interpolation, and backcast residual connections together induce the specialization of the additive predictions in different signal bands, reducing memory and computational time, thus improving the architecture’s parsimony and accuracy (Challu et al., 2023).